Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
Technical integration and coordination SASE1 Status and SASE3 Readiness
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 593     Entry time: 05 Mar 2025, 09:40
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 05.03.2025 
General Announcement:
- On the LIMP-2025 (accordingly to an e-mail sent by M. Doering on 24.02): reminder that groups are requested to give a feedback on the task list for the LIMP until tomorrow, 06.March.
- Next Friday UOM will have presentations from selected FD activities and jitter reports from instruments. 

CTRL:
- Nothing to report.

DA:
- Nothing to report.

EEE:
- Nothing to report.

VAC:
- Nothing to report.

XPD:
- Everything fine; busy being involved with FDs.

DET:
- Working on SPB replacing module for AGIPD; also busy with installation of AGIPD400K.

ITDM: 
- Nothing to report.

XRO:
- Nothing to report.

LAS:
- Not present.

UNSYS:
- Nothing to report.

SPB:
- Nothing apart of AGIPD intervention.

FXE:
- Nothing to report.

MID:
- Motor position issues while doing FD yesterday but experts are contacted.

HED:
- Nothing to report.

SCS:
- Jan Torben informs they are busy with installation; nothing to report.

SQS:
- Not present.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00